Solve the system: y = x2 + 5
6x - y = 3

Answer by Alan3354(69443)

y = x^2 + 5
6x - y = 3
-------
Sub for y in the 2nd eqn.
6x - x^2 - 5 = 3
---------
Solve for x, then for y